A satellite, moving in an elliptical orbit, is 306 km above Earth's surface at its farthest point and 168 km above at its closest point. Calculate (a) the semimajor axis (in km) and (b) the eccentricity of the orbit.
Answer #1

a = semimajot axis

e = eccentricity

Rf = farthest point = 306

Rc = closest point = 168

Rf = a(1 + e)                  eq-1

Rc = a(1 - e)                  eq-2

Rf/Rc = (1 + e) / (1 - e)

306 / 168 = (1 + e) / (1 - e)

e = 0.291

using eq-1

306 = a(1 + 0.291)

a = 237.03 km